OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x - 2y = 8. x intercept = (8, 0); y intercept = (0, 4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (-8, 0); y intercept = (0, -4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (anonymous):

To find the x intercept we make y = 0 x - 2y = 8 x - 2(0) = 8 x = 8 x intercept = (8,0) when y = 0 Do the same to find the y intercept. make x = 0 0 - 2y = 8 -2y = 8 can you finish it?

OpenStudy (anonymous):

-2/-2y=8/-2 y=-4 Is that correct??

OpenStudy (anonymous):

-8/-2 = 4 so the y intercept is (0,4) when x = 0. You were right but you forgot to do the negatives. A negative divided by a negative = positive.

OpenStudy (anonymous):

Oh okay, so the 8 was a negative?

OpenStudy (anonymous):

Oh wait I might of been wrong lol that is a +8 lol

OpenStudy (anonymous):

You are correct. -4 so the y intercept is (0,-4) it is late here and I am getting tired lol

OpenStudy (anonymous):

Me too. What time is it there where u live?

OpenStudy (anonymous):

11:00 pm

OpenStudy (anonymous):

Same time here too

OpenStudy (anonymous):

so the answer is the second one down.

OpenStudy (anonymous):

Don't forget the medal :-) I do like the shiny :-D

OpenStudy (anonymous):

OK Thanks!!

OpenStudy (whpalmer4):

yep, y-intercept is at (0,-4), x-intercept is at (8,0)